OpenStudy (anonymous):

A square ceiling has a diagonal of 24 ft. Shelton wants to put molding around the perimeter of the ceiling. How many feet of molding will he need? Round to the nearest foot.

OpenStudy (anonymous):



OpenStudy (anonymous):

A^2 + B^2 = C^2 is the formula you will use you know that C= 24ft, and that a and b must be the same number (because it is a square ceiling) 24^2 = A^2 + B^2 576 = A^2 + B^2 and since a and b must be the same, A^2 must equal half of 576 and B^2 must equal half of 576 so A^2 = 288 and B^2 = 288 which means that a= 16.97 and b=16.97 perimeter is 67.88 ft hope this helps!

OpenStudy (anonymous):

i will medal btw and i kind of need this quickly

OpenStudy (anonymous):

so it would be 17.0

OpenStudy (anonymous):

????

OpenStudy (anonymous):

perimeter is 67.88 ft, so that should be the amt of molding needed